OpenStudy (anonymous):

Decompose -6x/(x-6)(x+3) into partial fractions.

OpenStudy (anonymous):

What have you done so far?

OpenStudy (anonymous):

Since they basically give you the answer in the problem.

OpenStudy (anonymous):

I know x-6 and x+3 will be the denominators but I don't know how to get the numerator

OpenStudy (anonymous):

A/(x-6)+B/(x+3), write on the same denominator. After you've written them on the same denominator, compare what you have and what you want.

OpenStudy (anonymous):

-6x/[(x-6)(x+3)]=A/(x-6)+B/(x+3); as I said, write A and B on the same denominator, then you just compare with what you have and with what you need.

OpenStudy (anonymous):

For example, you have no constants in the left-hand-side.

OpenStudy (anonymous):

If there is a step you might not understand I'll explain it.

OpenStudy (anonymous):

It's crucial for learning that you solve problems yourself though -- once you have seen someone elses solution to an example, then you are deprived, forever, of the benefit that could have come from trying it yourself.

OpenStudy (john_es):

\[\frac{-6x}{(x-6)(x+3)}=\frac{A}{x-6}+\frac{B}{x+3}\] \[\frac{-6x}{(x-6)(x+3)}=\frac{Ax+3A+Bx-6B}{(x-6)(x+3)}=\frac{(A+B)x+3A-6B}{(x-6)(x+3)}\]Then \[A+B=-6\\ 3A-6B=0\Rightarrow A=-4, B=-2\] \[\frac{-6x}{(x-6)(x+3)}=\frac{-4}{x-6}+\frac{-2}{x+3}\] In case you still don't get the solution.

OpenStudy (anonymous):

You should refrain from giving out answers John_ES, please read the Code of Conduct on Openstudy.

OpenStudy (john_es):

I was waiting for the feedback of @ejune420, but I think that it should be best if she can check her solution. Anyways, the question was closed so I think there was no problem. Sorry for the trouble.

OpenStudy (anonymous):

I did not realize that the question had been closed; quite rude behavior from ejune420s part. Don't worry John_ES! :)

OpenStudy (john_es):

You were also right. It is better to help than provide the solution ;).

OpenStudy (anonymous):

lol rude behavior on my part? gtfo. my internet went down and didn't come back up until today. And did I ask for anyone to give me an answer? no? Didn't think so. I posted the question to get help, no where did I say someone give me an answer. Good try tho :) @Noliec

OpenStudy (john_es):

I believe it. In my case, openstudy crash my explorer/firefox/chrome randomly. Sometimes I even has written an answer that has been deleted by the crash.

OpenStudy (anonymous):

Yeah it crashes all the time on my google chrome. Thanks for not being an uppity jerk like that Noliec guy. You're awesome

OpenStudy (john_es):

Thanks, people should be more comprehensive with these problems that affect all of us.

OpenStudy (anonymous):

How did you manage to close the question with no internet connection? ;)

OpenStudy (anonymous):

I closed it once I got the answer, because I didn't need anyone else coming on helping, I already had you two on here to help me. That doesn't mean I was done coming to the question, I was trying to ask for an explanation when my internet went out. You've been a jerk from the start, so that's some rude behavior on your part. :)